# MOND as a transformation between non-inertial reference frames via Sciama’s interpretation of Mach’s Principle

### v4

@@ 1.2 Mach’s Principle and modified inertia
Schrödinger’s formulation appeared repeatedly in several works after him [22, 23, 24, 25, 26], but it was popularized by Dennis Sciama, who developed in 1953 a modified cosmological vector potential theory of inertia in which the inertial law arises as a side effect of gravity [27]. Sciama realized that in the analogous case of Maxwell’s equations applied to gravity, the rate of change of the vector potential leads to a term in the ‘gravitoelectric’ field that depends on the acceleration of an object relative to the rest of the masses of the universe. He showed that local inertial forces result from the gravitational induction of the universe, so that the dynamics of a rotating body would be affected by the large-scale distribution of the mass of the universe. Sciama postulated that “In the rest-frame of any body the total gravitational field at the body arising from all the other matter in the universe is zero”. The induced inertia decreased with $1/r$ with $r$ being the distance to the inertial body, so that the action of global matter dominates over the action of local matter, and it is not significantly modified by the acceleration with respect to local matter, leading to the illusion that inertia depends only on the body itself. The gravitational constant at any point was determined by the total potential of the distribution of matter of the universe, taking the form of $G\sim c^{2}/(\Phi_{u}+\phi)$ with $\Phi_{u}=M_{u}/R_{u}$ and local $\phi=M/r$ (assuming a linear superposition) which reduces to $G=c^{2}/(M_{u}/R_{u})$ in most cases (when locally $\Phi_{u}>>\phi$, since only near neutron stars or black holes $\Phi_{u}\sim\phi$) with $M_{u}$ and $R_{u}$ the mass and radius of the observable universe, causally connected by the speed of light in the past light-cone of the point considered to evaluate $G$, for the equivalence of inertial and gravitational masses to hold. In this way, local phenomena were strongly coupled to global properties of the universe. Sciama knew about the Hubble expansion of the universe, but his model predicted an almost flat observable universe in which critical cosmic matter density is reached for $G\Phi_{u}/c^{2}=1$ to hold, even though the flatness of the universe was not observed at his time.
@@
By pure dimensional analysis, the Reissner-Schrödinger- Sciama’s relationship (henceforth Sciama’s relationship) is the only possible derivation of the units of measurement of the gravitational constant through a mass, a distance, and the speed of light. If Sciama’s relationship is derived from a more fundamental theory, it can be expected to also contain integers and mathematical constants, but these will be omitted for simplicity, since it does not significantly affect the resulting orders of magnitude of the comparison between $M_{u}$ and $R_{u}$.
@@
Robert H. Dicke and Carl H. Brans attempted to introduce Sciama’s relationship in a relativistic scalar-tensor theory of gravitation in 1961 [28], inspired by Mach’s principle, by interpreting the scalar field as an advanced wave integral over all matter. One interpretation of Brans-Dicke theory consists in the variation of the inertial mass with the scalar field if the gravitational constant is constant, so that Einstein’s field equations are formally valid, but the most well-known interpretation allows the gravitational constant to be variable with position and time. The difference in predictions between the Brans-Dicke theory and GR has been locally tested in the solar system near the Sun through the Shapiro time delay effect measured by the Cassini experiment, showing that the Brans-Dicke dimensionless coupling factor of the scalar field to gravity, which differentiates between the theory and GR, must be so high that Brans-Dicke theory is almost indistinguishable from GR. All of these tests have been performed in the high-acceleration regime of the solar system.
@@
Hans-Jürgen Treder developed a model similar to that of Schrödinger in a more complete inertia-free mechanics model [23] based on the Riemann potential as a velocity-dependent gravitational potential (instead of the Weber potential of Reissner and Schrödinger). This model implements the idea of inertia having a gravitational origin without Schrödinger’s and Sciama’s anisotropic inertial mass (which is ruled out by observations, such as the Hughes–Drever experiments) with Sciama’s $G\sim c^{2}/(\Phi_{u}+\phi)$, in agreement with Mach’s principle. It is worth noting that Treder was able to implement Mach’s principle maintaining a scalar inertial mass instead of a tensorial one, exhibiting no preferred direction and thus avoiding inertial anisotropy. In this non-relativistic linear model, the weak equivalence principle is derived and not axiomatically postulated as in GR, the weakness of the gravitational constant is explained, and inertial mass is also derived from the model.
@@
Sciama’s relationship with today’s measurements of the current cosmological model is satisfied only in terms of orders of magnitude, because the Hubble tension impedes a precise calculation. It yields the orders of magnitude of the observed gravitational constant when considering not the baryonic and dark matter content for $M_{u}$, but the total energy content (since all forms of energy must gravitate according to GR), which roughly corresponds to the critical density of the universe as measured to be almost flat. As Milgrom stated in one of his conferences, “the only system that is strongly general relativistic and in the MOND regime is the Universe at large”. However, the critical density is calculated with the assumption of a constant gravitational constant itself and considering the dark matter content, which could be reduced through an effective correction to Newtonian laws, as will be proposed in the next section.
@@
Few authors have attempted to develop a MOND for galaxy rotation curves based on Mach’s principle [29], or relate Milgrom’s MOND to Sciama’s relationship and Mach’s principle [30, 31, 32, 33, 34, 35, 36]. The most relevant ideas in the literature relating these topics are summarized onwards.
@@
Alexander Unzicker explored galaxy dynamics in relation to Mach’s Principle and the rotating bucket of water experiment, and proposed a basis for a MOND to solve flat rotation curves without considering Milgrom’s MOND [29]. He proposes the following thought experiment. According to Mach, the system of two distant masses rotating around their center of mass in equilibrium of gravitational and centrifugal forces in an empty universe, since rotation between them is undefined due to absence of absolute space, must be equivalent to the system of those two distant masses not rotating and without gravity (only radial or relative velocities are measurable, and tangential velocities are impossible to measure instantaneously, as they are indistinguishable from a rotation of the coordinate system of the observer). The maximum possible angular velocity is limited by $w_{max}=c/r$ so that the maximum tangential speed is $v=c$ for one of the masses rotating around the other one, since the speed of light cannot be surpassed in any case. This luminal rotational speed in absence of a background universe from Unzicker’s thought experiment is a feature of introducing Sciama’s relationship in Newton’s law of gravity, in which the Newtonian $v=\sqrt{GM/r}$ changes to $v=c\sqrt{(M/r)/(M_{u}/R_{u})}$ after introducing $G=c^{2}/(M_{u}/R_{u})$ and without a background universe $M_{u}\rightarrow M$ and $R_{u}\rightarrow r$, velocities are $v=c$. In other words, Sciama’s relationship has the equivalent effect of decreasing inertia when decreasing $M_{u}$.
@@
Jaume Gine explored a Machian interpretation of the modified second law of motion for the simple interpolating function of Milgrom’s MOND based on the accelerated expansion of the universe [33], in which $a_{0}=GM_{u}/{R_{u}}^{2}=c^{2}/R_{u}$ through Sciama’s relationship is the acceleration that an experimental body feels induced by the rest of the matter of the universe in its inertial frame of reference, and the value of acceleration for which inertial and gravitational masses of a body can differ. Thus, the equivalence principle between inertial and gravitational masses is broken for accelerations smaller than $a_{0}$. In a following paper [34], Gine attempted to derive a phenomenological version of Milgrom’s MOND modification to Newton’s second law. He considers $a_{0}$ to be the acceleration at which the edge of the universe at distance $R_{u}$ goes away from a considered central inertial point due to the expansion of the universe. By relativizing acceleration considering the distant universe at rest, he derives a new interpolating function similar to Milgrom’s MOND simple interpolating function.
@@
A complete FundaMOND theory remains only a faint hope until the origin of Milgrom’s MOND is correctly identified. Even though interesting ideas have been put forward relating MOND to Mach’s principle, an exact equivalence based on the Machian consequence that the inertia of a body should decrease when masses are removed from its neighborhood (at the deep MOND regime) will be presented, which should arise as an approximation in any non-linear theory of modified inertia or gravity which attempts to explain galaxy rotation curves. The motivation behind this interpretation of Milgrom’s MOND is that MOND can be reformulated with the same parameters of the mass and radius of the observable and causally connected universe that the Machian models of modified inertia contain.
2 Milgrom’s MOND from Mach’s Principle
@@
Starting from the well-known form of Milgrom’s MOND standard or simple interpolating function correction to Newtonian gravity or to inertia, the ‘true’ gravitational acceleration $g$ or ‘true’ second law of inertia is
@@
with $a_{0}$ being Milgrom’s MOND acceleration scale constant, $g_{N}=GM/r^{2}$ the Newtonian gravitational acceleration, $M$ the active gravitational mass, $m_{i}$ the inertial mass, and $r$ the distance between the centers of mass of the active and passive gravitational masses. For the motion of a test particle in a gravitational field in the low acceleration or deep MOND regime, $\mu\left(a/a_{0}\right)\rightarrow a/a_{0}$ and $g=\sqrt{a_{0}g_{N}}$, so that with $g=v^{2}/r$, $v=\sqrt[4]{GMa_{0}}$ for achieving flat galaxy rotation curves in agreement with the Tully-Fisher law in a remarkable simple way with a single new free parameter. For the high acceleration regime, $\mu\left(a/a_{0}\right)\rightarrow 1$ and Newtonian laws are recovered.
Milgrom’s MOND acceleration scale constant $a_{0}$ is usually interpreted as a fundamental constant in MOND theory, but possibly related to the cosmological parameter of the cosmological constant $a_{0}\sim c^{2}\sqrt{\Lambda}$ or the energy density of the vacuum $\rho_{vac}$ with $\Lambda=8\pi G\rho_{vac}/c^{2}$ according to GR, which coincides in orders of magnitude (a better match can be done with $a_{0}=c^{2}\sqrt{\Lambda/3}$, although integers and Pi are onwards omitted for simplicity). This is due to Milgrom’s wish to define absolute acceleration with respect to a smooth vacuum energy frame in MOND as modified inertia. But Mach’s principle implies that acceleration should be defined with respect to the rest of the masses of the universe. Milgrom’s reluctance to explore a Machian version could be attributed to the widespread false claim that Mach’s principle necessarily implies anisotropy of inertia (due to non-uniform mass distributions), strongly constrained by the Hughes-Drever experiments. As shown by Treder, as long as inertial mass is a scalar, it exhibits no preferred direction, and no anisotropy of inertia. In order to express $a_{0}$ in terms of the mass and radius of the universe, $\rho_{vac}$ is substituted for the matter density of the universe $\rho_{vac}\sim\rho_{u}\sim M_{u}/{R_{u}}^{3}$ with $M_{u}$ the causally connected mass to the local system of study through the radius $R_{u}$ of the observable universe (or particle horizon radius), a relationship which also coincides in orders of magnitude. Newton’s gravitational constant $G$ is substituted for Sciama’s interpretation of Mach’s principle $G\sim c^{2}/(M_{u}/R_{u})$, so that $a_{0}\sim GM_{u}/{R_{u}}^{2}\sim c^{2}/R_{u}$, and by reverse engineering, $v\sim c\sqrt[4]{M/M_{u}}$ and $g\sim c^{2}\sqrt{M/M_{u}}/r$ in the equivalent deep MOND regime. The interpolating function is now scale-invariant in mass and length, resulting in a Machian MOND formulation
$g=\frac{c^{2}}{\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)\ \mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)}\frac{M}{r^{2}}\ \ or\ \ \frac{M}{r^{2}}=\frac{\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)}{c^{2}}\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)g$ (3)
$\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)=\frac{1}{\sqrt{1+\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}}}\ \ or\ \ \mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)=\frac{1}{{1+\sqrt{\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}}}}$ (4)
@@
in which for most cases where local potentials are small, $\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)\sim\left(\frac{M_{u}}{R_{u}}\right)$.
It is trivial that, since only equivalent substitutions for $G$ and $a_{0}$ according to Sciama’s relationship based on Mach’s principle which yield around the same values are done (omitting integers and mathematical constants), and $M$ and $r$ take the same meaning as in Milgrom’s MOND, the resulting corrections (3) and (4) for galaxy rotation curves are equivalent to Milgrom’s MOND. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}<<1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow 1$ and Newtonian laws are restored with $v=c\sqrt{({M/r})/({M_{u}/R_{u}})}$. For $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$, $\mu\left(\sqrt{\frac{M/{r}^{2}}{M_{u}/{R_{u}}^{2}}}\right)\rightarrow\sqrt{\frac{M/r^{2}}{M_{u}/{R_{u}}^{2}}}$, $v=c\sqrt[4]{M/M_{u}}$, and the speed of rotating stars around galaxies far away from their galactic centers is determined by the relative active gravitational masses between the local and the global system.
It is also trivial that the denominator $M_{u}/{R_{u}}^{2}$ in (4) is the total field intensity (the gravitational field intensity without the gravitational constant $G$) of all masses of the universe (omitting dimensionless factors for simplicity), taking the same physical meaning as $M/{r}^{2}$ in the numerator, which is the field intensity of the masses of the local system at the point in which the correction is applied.
Thus, a transformation (4) can be identified between the non-inertial reference frame of a local system with a mass parameter $M$ and distance $r$, and a global reference frame of observable mass $M_{u}$ and size $R_{u}$ of the universe. The variable changed by this transformation is either the gravitational force (equivalent to a variable gravitational constant) or the inertia of a body. Note that in (1), the interpolating function depends on true acceleration, while in Machian MOND it depends on field intensity. In particular, as modified inertia, it depends on the inertial mass, the local mass, and the mass of the universe, showing that inertia arises from the overall mass distribution à la Mach. It also makes use of the same parameters $M_{u}$ and $R_{u}$ as the modified inertia theories of Sciama and Treder.
By decreasing the mass of the universe within the observable radius, the transformation (4) decreases the gravitational force (or increases inertia) when $\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}>>1$. But, considering also Sciama’s relationship for the gravitational constant, the gravitational force increases (having the same observational consequence as considering Sciama’s relationship alone, the decrease of inertia), and local velocities increase. The formal limit where $a_{0}\rightarrow 0$ in which Milgrom’s MOND approaches Newton’s laws does not hold, since the mass and size of the universe (for which $a_{0}$ is effectively replaced) always exist. The mass of the universe can never be zero in the system under study, it can only be small as the very same mass of the system itself. Decreasing the mass of the universe until the absence of any background ($M_{u}\rightarrow M,R_{u}\rightarrow r$), the transformation (4) reduces to some simple number $K$ which can be set to unity with the appropriate choice of integers in the formulation of the Machian transformation and in Sciama’s relationship (a simple way to obtain $K=1$ is considering Schrödinger’s original definition of $G$ and multiplying the relational field term of the transformation by $3$), and velocities approach $v\rightarrow Kc$. This is already a Machian feature of Sciama’s relationship by which in absence of a background universe, rotation is undefined, and tangential speed can take any value up to the limit of the speed of light. The proposed transformation respects this result and also respects the scale invariance in mass, length, and time that Sciama’s relationship introduces in the Newtonian regime, while classical Newtonian gravity is scale-dependent. Milgrom’s MOND is only length and time scale-invariant when $a_{0}\rightarrow\infty$ fixing $Ga_{0}$ to achieve constant velocities independent of radius at the flat part of the galaxy rotation curves. Machian MOND is thus scale-invariant in mass, length, and time $(m,r,t)\rightarrow(\lambda m,\lambda r,\lambda t)$ at all regimes. Scaling the mass or size of both the universe and the local system always results in the same observational velocities.

### v5

@@ 1.2 Mach’s Principle and modified inertia
Schrödinger’s formulation appeared repeatedly in several works after him [Whitrow1946, Treder1972, WhitrowRandall1951, Dicke1961, Assis1999], but it was popularized by Dennis Sciama, who developed in 1953 a modified cosmological vector potential theory of inertia in which the inertial law arises as a side effect of gravity [Sciama1953]. Sciama realized that in the analogous case of Maxwell’s equations applied to gravity, the rate of change of the vector potential leads to a term in the ‘gravitoelectric’ field that depends on the acceleration of an object relative to the rest of the masses of the universe. He showed that local inertial forces result from the gravitational induction of the universe, so that the dynamics of a rotating body would be affected by the large-scale distribution of the mass of the universe. Sciama postulated that “In the rest-frame of any body the total gravitational field at the body arising from all the other matter in the universe is zero”. The induced inertia decreased with $1/r$ with $r$ being the distance to the inertial body, so that the action of global matter dominates over the action of local matter, and it is not significantly modified by the acceleration with respect to local matter, leading to the illusion that inertia depends only on the body itself. The gravitational constant at any point was determined by the total potential of the distribution of matter of the universe, taking the form of $G\sim c^{2}/(\Phi_{u}+\varphi)$ with $\Phi_{u}=M_{u}/R_{u}$ and local $\varphi=M/r$ (assuming a linear superposition) which reduces to $G=c^{2}/(M_{u}/R_{u})$ in most cases (when locally $\Phi_{u}>>\varphi$, since only near neutron stars or black holes $\varphi\sim\Phi_{u}$) with $M_{u}$ and $R_{u}$ the mass and radius of the observable universe, causally connected by the speed of light in the past light-cone of the point considered to evaluate $G$, for the equivalence of inertial and gravitational masses to hold. In this way, local phenomena were strongly coupled to global properties of the universe. Sciama knew about the Hubble expansion of the universe, but his model predicted an almost flat observable universe in which critical cosmic matter density is reached for $G\Phi_{u}/c^{2}=1$ to hold, even though the flatness of the universe was not observed at his time.
@@
By pure dimensional analysis, the Reissner-Schrödinger-Sciama’s relationship (also known as the Hofmann-Reissner-Schrödinger relationship, henceforth Sciama’s relationship) is the only possible derivation of the units of measurement of the gravitational constant through a mass, a distance, and the speed of light. If Sciama’s relationship is derived from a more fundamental theory, it can be expected to also contain integers and mathematical constants, but these will be omitted for simplicity, since it does not significantly affect the resulting orders of magnitude of the comparison between $M_{u}$ and $R_{u}$. From this relationship, it became evident that the inability to explain the gravitational constant in both Newton’s and Einstein’s theories is related to the inability to explain inertia à la Mach and to consider the universe at large.
@@
Carl H. Brans introduced a thought experiment considering the ratio between the inertial mass of a body and the active gravitational mass following Mach’s principle in a static universe consisting only of a mass shell of radius $R_{u}$ and inertial mass $M_{u}$ (these will be the mass and size of the universe), together with a relatively small body of inertial mass $M$ and a test particle. He expected the acceleration of the test particle to be independent from the mass of the particle itself (from the equivalence principle), to depend on the mass $M$, and because of Machian arguments about the inertial properties of the test body being determined by the distribution of matter in the universe and not intrinsic of the body, conceivably to also depend on $M_{u}$ and $R_{u}$. Thus, $a=f(M_{u},R_{u})M/r^{2}$ and he related $f$ to the gravitational constant [Brans1962]. Together with Robert Dicke, they attempted to introduce Sciama’s relationship $G=c^{2}/(M_{u}/R_{u})$ in a relativistic scalar-tensor theory of gravitation in 1961 [Dicke1961], inspired by Mach’s principle, by interpreting the scalar field as an advanced wave integral over all matter. One interpretation of Brans-Dicke theory consists in the variation of the inertial mass with the scalar field if the gravitational constant is constant, so that Einstein’s field equations are formally valid, but the most well-known interpretation allows the gravitational constant to be variable with position and time. The difference in predictions between the Brans-Dicke theory and GR has been locally tested in the solar system near the Sun through the Shapiro time delay effect measured by the Cassini experiment, showing that the Brans-Dicke dimensionless coupling factor of the scalar field to gravity, which differentiates between the theory and GR, must be so high that Brans-Dicke theory is almost indistinguishable from GR. All of these tests have been performed in the high-acceleration regime of the solar system.
@@
Hans-Jürgen Treder developed a model similar to that of Schrödinger in a more complete inertia-free mechanics model [Treder1972] based on the Riemann potential as a velocity-dependent gravitational potential (instead of the Weber potential of Reissner and Schrödinger). This model implements the idea of inertia having a gravitational origin without Schrödinger’s anisotropic inertial mass (which is ruled out by observations, such as the Hughes–Drever experiments) with Sciama’s $G\sim c^{2}/(\Phi_{u}+\varphi)$, in agreement with Mach’s principle. It is worth noting that Treder was able to implement Mach’s principle maintaining a scalar inertial mass instead of a tensorial one, exhibiting no preferred direction and thus avoiding inertial anisotropy. In this non-relativistic linear model, the weak equivalence principle can be derived and not axiomatically postulated as in GR, the weakness of the gravitational constant is explained, and inertial mass can also derived from the model.
@@
Sciama’s relationship with today’s measurements of the current cosmological model is satisfied only in terms of orders of magnitude, because the Hubble tension impedes a precise calculation. It yields the orders of magnitude of the observed gravitational constant when considering not the baryonic and dark matter content for $M_{u}$, but the total energy content (since all forms of energy must gravitate according to GR), which roughly corresponds to the critical density of the universe as measured to be almost flat. As Milgrom stated in one of his conferences, “the only system that is strongly general relativistic and in the MOND regime is the Universe at large”. However, the critical density is calculated with the assumption of a constant gravitational constant itself and considering dark matter, which would not contribute if a correction to the laws of gravity and inertia accounts for its effects, as it will be proposed in the next section.
@@
Few authors have attempted to develop a MOND for galaxy rotation curves based on Mach’s principle [Unzicker2003], or relate Milgrom’s MOND to Sciama’s relationship and Mach’s principle [Darabi2010, Corda2016, Anderson2011, Gine2009, Gine2012, Abreu2014, Corda2024]. The most relevant ideas in the literature relating these topics are summarized onwards.
@@
Alexander Unzicker explored galaxy dynamics in relation to Mach’s Principle and the rotating bucket of water experiment, and proposed a basis for a MOND to solve flat rotation curves without considering Milgrom’s MOND [Unzicker2003]. He draws attention to the following thought experiment. According to Mach, the system of two distant masses rotating around their center of mass in equilibrium of gravitational and centrifugal forces in an empty universe, since rotation between them is undefined due to absence of absolute space, must be equivalent to the system of those two distant masses not rotating and without gravity (only radial or relative velocities are measurable, and tangential velocities are impossible to measure instantaneously, as they are indistinguishable from a rotation of the coordinate system of the observer). The maximum possible angular velocity is limited by $w_{max}=c/r$ so that the maximum tangential speed is $v=c$ for one of the masses rotating around the other one, since the speed of light cannot be surpassed in any case. This luminal rotational speed in absence of a background universe from Unzicker’s thought experiment is a feature of introducing Sciama’s relationship in Newton’s law of gravity, in which the Newtonian $v=\sqrt{GM/r}$ changes to $v=c\sqrt{(M/r)/(M_{u}/R_{u})}$ after introducing $G=c^{2}/(M_{u}/R_{u})$ and without a background universe $M_{u}\rightarrow M$ and $R_{u}\rightarrow r$, velocities are $v=c$. In other words, Sciama’s relationship has the equivalent effect of decreasing inertia when decreasing $M_{u}$.
@@
Jaume Gine explored a Machian interpretation of the modified second law of motion for the simple interpolating function of Milgrom’s MOND based on the accelerated expansion of the universe [Gine2009], in which $a_{0}=GM_{u}/{R_{u}}^{2}=c^{2}/R_{u}$ is the acceleration that an experimental body feels induced by the rest of the matter of the universe in its inertial frame of reference, and the value of acceleration for which inertial and gravitational masses of a body can differ. Thus, the equivalence principle between inertial and gravitational masses is broken for accelerations smaller than $a_{0}$. In a following paper [Gine2012], Gine attempted to derive a phenomenological version of Milgrom’s MOND modification to Newton’s second law. He considers $a_{0}$ to be the acceleration at which the edge of the universe at distance $R_{u}$ goes away from a considered central inertial point due to the expansion of the universe. By relativizing acceleration considering the distant universe at rest, he derives a new interpolating function similar to Milgrom’s MOND simple interpolating function.
@@
A complete FundaMOND theory remains only a faint hope until the origin of Milgrom’s MOND is correctly identified. Even though interesting ideas have been put forward relating MOND to Mach’s principle, a formulation based on Einstein’s Machian consequence not satisfied by GR, by which the inertia of a body should decrease when masses are removed from its neighborhood, will be presented in the next section. Such formulation should arise as an approximation in any non-linear Machian theory of modified inertia or gravity which attempts to explain galaxy rotation curves. The motivation behind this interpretation of Milgrom’s MOND is that MOND can be reformulated with the same parameters of the mass and radius of the observable and causally connected universe that the Machian models of modified inertia are built on.
2 Machian MOND
@@
Starting from the well-known form of Milgrom’s MOND standard or simple interpolating function correction to Newtonian gravity or to inertia, the ‘true’ gravitational acceleration $g$ or ‘true’ second law of motion is
@@
with $a_{0}$ being Milgrom’s MOND acceleration scale constant, $g_{N}=GM/r^{2}$ the Newtonian gravitational acceleration, $M$ the active gravitational mass, $m_{i}$ the inertial mass, and $r$ the distance between the centers of mass of the active and passive gravitational masses. For the motion of a test particle in a gravitational field in the low acceleration or deep MOND regime, $\mu\left(a/a_{0}\right)\rightarrow a/a_{0}$ and $g=\sqrt{a_{0}g_{N}}$, so that with $g=v^{2}/r$, $v=\sqrt[4]{GMa_{0}}$ to achieve flat galaxy rotation curves in agreement with the TF law in a remarkable simple way with a single new constant $a_{0}$. For the high acceleration regime, $\mu\left(a/a_{0}\right)\rightarrow 1$ and Newtonian laws are recovered. One can also define an equivalent interpolating function $g=\nu(a_{0}/a_{N})a_{N}$ converting the Newtonian gravitational field into a Milgromian field, with $\nu(a_{0}/a_{N})=(1/2)^{1/n}(1+\sqrt{1+4(a_{0}/a_{N})^{n}})^{1/n}$ (with $n=1$ for the simple and $n=2$ for the standard interpolating functions).
Milgrom’s MOND acceleration scale constant $a_{0}$ is usually interpreted as a fundamental constant in MOND, but possibly related to the Hubble parameter $a_{0}\sim cH_{0}$ or the cosmological constant $a_{0}\sim c^{2}\sqrt{\Lambda}\sim c^{2}/R_{u}$ as the energy density of the vacuum $\rho_{vac}$ with $\Lambda=8\pi G\rho_{vac}/c^{2}$ according to GR, which coincides in orders of magnitude (a better match can be done with $a_{0}=c^{2}\sqrt{\Lambda/3}$, although integers and Pi are onwards omitted for simplicity). This is due to Milgrom’s wish to define absolute acceleration with respect to a smooth quantum vacuum energy frame in MOND as modified inertia. However, the cosmological constant is only measured indirectly, and the connection between dark energy and a non-zero energy density of the quantum vacuum remains, at best, an optimistic hypothesis. In contrast, Mach’s principle implies that acceleration should be defined with respect to the mass distribution of the universe, and the mass and size of the universe are based on direct, albeit approximate measurements.
In order to express $a_{0}$ in terms of the mass and radius of the universe, the cosmic coincidence implies $\rho_{vac}\sim\rho_{u}\sim M_{u}/{R_{u}}^{3}$ with $M_{u}$ the causally connected mass to the local system of study through the radius $R_{u}$ of the observable universe (or particle horizon radius). Newton’s gravitational constant $G$ is substituted for Sciama’s interpretation of Mach’s principle $G\sim c^{2}/(M_{u}/R_{u})$, so that $a_{0}\sim GM_{u}/{R_{u}}^{2}$, and by reverse engineering, $v\sim c\sqrt[4]{M/M_{u}}$ and $g\sim c^{2}\sqrt{M/M_{u}}/r$ in the equivalent deep MOND regime. By pure dimensional analysis, $a_{0}\sim GM_{u}/{R_{u}}^{2}$ is also the only possible equation to obtain an acceleration from $M_{u}$, $R_{u}$ and other fundamental constants. The interpolating function is now invariant under global rescaling of mass and length, resulting in a Machian MOND formulation
$g=\frac{c^{2}}{\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)\ \mu\left(\frac{M_{u}/{R_{u}}^{2}}{M/{r}^{2}}\right)}\frac{M}{r^{2}}\ \ or\ \ \frac{M}{r^{2}}=\frac{\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)}{c^{2}}\mu\left(\frac{M_{u}/{R_{u}}^{2}}{M/{r}^{2}}\right)g$ (3)
$\mu\left(\frac{M_{u}/{R_{u}}^{2}}{M/{r}^{2}}\right)=\frac{1}{\sqrt{1+\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}}}\ \ or\ \ \mu\left(\frac{M_{u}/{R_{u}}^{2}}{M/{r}^{2}}\right)=\frac{1}{{1+\sqrt{\frac{M_{u}/{R_{u}}^{2}}{M/r^{2}}}}}$ (4)
@@
in which for most cases where local potentials are small, $\left(\frac{M_{u}}{R_{u}}+\frac{M}{r}\right)\sim\left(\frac{M_{u}}{R_{u}}\right)$. The first Machian interpolating function is the best fit to Milgrom’s MOND $\nu(a_{0}/a_{N})$.